determine whether the line through the pair of points A, B is
parallel or perpendicular to the line through the pair C, D.
3.) A(1,5), B(-7,1) and C(8,-2), D(4,6)
4.) A(6,2), B(8,6) and C(4,8), D(2,4)
1
Expert's answer
2020-05-05T20:13:46-0400

3) To find an angle, we should find slope "k_1"

for AB line and slope "k_2" for CD line.

If we will choose equation of line AB as "y=k_1\cdot x + b_1" so, we will have a system: "\begin{cases}5=k_1\cdot1+b_1
\\1=k_1\cdot(-7)+b_1\end{cases}"From it, we could find that "k_1=0.5"

Like that, CD line has equation "y=k_2\cdot x+b_2" and "k_2=-2" .

As "k_1\cdot k_2 = -1" it means that, these lines are perpendicular. The answer: AB and CD are perpendicular.



4) As the question above, we could find that "k_1=2,\space k_2=2" . As, "k_1 = k_2", AB and CD lines have equal slopes, more clearly, they are parallel lines. Answer: AB and CD lines are parallel